OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 6 and the last term is 1,000,842?

OpenStudy (anonymous):

a. 138,073 b. 117,648 c. 188,851 d. 157,962

OpenStudy (anonymous):

First you have to do: \[a_{6} = a_{1}r^{6 - 1}\]\[1000842 = 6r^{5}\]Can you solve for r now?

OpenStudy (anonymous):

What do I do with the 5?

OpenStudy (anonymous):

Well, did you first divide by 6?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[r = \sqrt[5]{166807}\]

OpenStudy (anonymous):

After you get r, then plug it into this formula. \[S_{n} = \frac{a_{1}(1 - r^{n})}{1 - r}\]\[S_{6} = \frac{6(1 - r^{6})}{1 - r}\]

OpenStudy (anonymous):

Note: You need a scientific caluculator or use a web calculator.